OpenStudy (anonymous):

Two of the interior angles of a triangle are equal. The third interior angle is 99.4 degrees. What are the other two interior angles? theta=?degress

OpenStudy (anonymous):

What is the sum of the 3 angles of a triangle?

OpenStudy (anonymous):

Do you remember that the sum is 180 degrees?

OpenStudy (anonymous):

Let x be the measure of the two equal angles, then x+ x +99.4=180 Solve for x

OpenStudy (anonymous):

\[ 2 x =180-99.4\\ 2 x= 80.6\\ x= 40.3 \]

OpenStudy (anonymous):

thank you!